If a cube of side 5 cm has a charge of 6 microcoulombs , then the surface charge density is

Grade:12

2 Answers

hero
6 years ago
surface charge density is equal to total charge on surface divided by the surface area.
surface area of cubeis 6a2. a= 5 cm =0.05 m .
area= 6 ×0.052=0.015m2
surface charge density= charge/area
                                 =6 ×10-6/0.015
                                 =4 × 10-4 coulomb/m2.
hope it helps
 
ankit singh
3 years ago

 

(a) Side =5cm=0.05m  area =6a2=6×(0.05)2
charge =6μC
The surface density =6a2q
                                  =6×(0.05)26×106
                                  =1005×1005106
                                  =400×106
                                  =4×104C/m2